Solve the following spherical triangle using Law of Sine and Cosine.

2. A= 82°, b=100°, c=78° Find: B


1
Expert's answer
2022-02-09T13:33:00-0500

Using the Law of Cosine, we can find the side a.


"cos (a) = cos(b) * cos (c)+sin(b)*sin(c)*cos(A)"

"cos(a)= cos(100\u00b0) * cos (78\u00b0)+sin(100\u00b0)*sin(78\u00b0)*cos(82\u00b0)"

"cos(a)= -0.74 * 0.208+0.985*0.978*0.139="

"=-0,154+0.134=0.02"

"a=arccos(0.02)=88.85\u00b0"

And then, using the Law of Sines, we can find B.


"sin(a)\/sin(A)=sin(b)\/sin(B)"

"sin(B)=sin(b)*sin(A)\/sin(a)"

"sin(B)=sin(100\u00b0)*sin(82\u00b0)\/sin(88.85\u00b0)"

"sin(B)=0.9848*0.9903\/0.9998=0.9754"

"B=arcsin(0.9754)=77.27\u00b0"

Answer: 77.27°